02-03 October 2008 Alignment of the New WFS Camera

Observers: D. Thompson and J. Hill in Tucson
Instrument Support: (Working hard on the platform above the primary mirror) T. Shih, and earlier J. Little
Software Support: N. Cushing in Tucson
Telescope Operator: A. Ceranski

X-Y adjustment of the new WFS camera with the calibration lamp in place.

Times below in UT, even though we aren't on the sky.

We are now at the 5th step of Jesper's emailed procedure for replacing the WFS camera. On the old camera, the on-axis spots were centered at 395,260 (image wfsc 15 from two nights ago). They moved to 396,286 when the probe is on an off-axis star (gs=82 for image wfsc 32 from two nights ago). We are aiming to get the spots slighly below center in the on-axis position so that we don't lose the top row of spots as the probe moves off-axis.

03:10 Tim's "Out" adjustment moves the spots up on the DS9 display. Highest useable row is at Y=326. Images are wfsc 110 and nearby.

03:14 Lowest useable row is Y=142. e.g. image wfsc 153 Therefore the center row is Y=234.

Our target for the center of the calibration pupil is Y=234-21=213.

The adjusting screws are 0.5 mm pitch. The pitch of the spots in the S-H array is 13 CCD pixels. Tim is impressively able to make 2 pixel adjustments by making fractional turns of the screws.

03:42 Tim's "Left" moves spots left. Leftmost useable column is X=308.

Rightmost useable column appears to be X=480, although one of the mounting screws seems to be stopping us from moving the pupil spots any farther to the right. Based on dome flats taken later (wfsc 1121, 1122) it looks like column 15 of the lenslets is at X=493.

3:58 We've concluded that the world is pretty happy with the calibration lamp pupil centered at X=388 Y=217. This leaves us missing the bottom row of the calibration lamp pupil because it is larger than the telescope pupil.

04:03 All screws tight with the adjustment bracket removed. Final position is shown in wfsc 1116.

05:00 We are running the AGw unit, making sure that everything seems to be working with the dome closed.

05:30 GCS hangs while we are taking a 1 ms WFS image. We did a stopAGW, killed GCS, startAGW, and restarted GCS.

05:45 Dave does stowAGW and kills the oacsever windows we have displayed downtown so that Aaron can restart them on the to-station.

I've copied some of the images we've taken from /tmp on lbmu107 to /Repository/AGW_Data/081003w_gcs/.

No observing for the rest of tonight since: a) work platform is still in place, b) they are purposefully letting the HBS oil settle to get rid of some water that got in it, c) the weather is mostly cloudy.

Tomorrow daytime

We believe that we are all finished with this WFS camera alignment procedure. Tim has removed the calibration lamp and buttoned everything up. So the AGw is ready to go. Jon R. and friends should remove the work platform in the morning. Then we need to put the swing arms back in LUCIFER configuration and rebalance the telescope.

Tomrrow or whenever we get on a star

Need to adjust GCS WFScam_hotspot_x and ....._y in /lbt/tcs/current/etc/GCS/AIP_L.cfg to define the position of the spot array on the WFS image. Old values were X=395 Y=261. New values based on the calibration lamp will be approximately X=388 Y=217.

Andrew reports that Doug's IDL program will give us the option of resetting the center position so we can work out the new numbers.

-- JohnHill - 03 Oct 2008
Topic revision: r2 - 03 Oct 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback